OpenStudy (anonymous):

What is the solution set of 2x(x - 1) = 3

OpenStudy (anonymous):

Complete the square: expand: 2x^2 - 2x = 3 divide by 2 x^2 - x = 3/2 add the square of half of the coefficient on x to both sides. x^2 - x + 1/4 = 3/2 + 1/4 x^2 - x + 1/4 = 7/4 factor: (x - 1/2)^2 = 7/4 square root: |x - 1/2| = sqrt(7)/2 x -1/2 = sqrt(7)/2 or x - 1/2 = -sqrt(7)/2 x = sqrt(7)/2 + 1/2 or x = 1/2 - sqrt(7/2) or: x = 1/2(1 +/- sqrt(7))

jhonyy9 (jhonyy9):

2x*2 -2x -3 =0 2 +/- radical ( 4 + 24) 2+/- radical 28 2+/- 2radical 7 x_1,_2 = ------------------ = ------------- = ------------- = 4 4 4 2(1+/- radical 7) 1+/- radical 7 = -------------- = ------------ 4 2